Each of the following lines has a given length. Determine the engineer's scale used in each line. Express in the form 1'' = x' or similar.

1) 800' ــــــــــــــــــــــ scale =?

2)126 miles ــــــــــــــــــ scale =?

3) 2.1' ــــــــــــــــــــــ scale =?

4) 23,700' ــــــــــــــــــــــ scale =?

Each of the following lines has a given length. Determine the architect's scale used in each line. Express your answer in the form x''/y = 1':

1) 14'- 3'' ــــــــــــــــــــــ scale =?

2) 2'-1 1/4'' ــــــــــــــــــــــ scale =?

3) 8'- 5'' ــــــــــــــــــــــ scale =?

4) 3'- 11 1/2'' ــــــــــــــــــــــ scale =?
Answer #1

Simply measure the lines using the Engineer's scale and whatever value is coming to take its ratio with the given length of the line.

1) Engineer's scale reading = 0.76 inches, actual distance on ground = 800' = 800*12 inches

so scale = 0.76 : 9600 = 1:12631.578, this means that 1 inch on the map equals 12631.57 on ground.

2)  Given line measures equal 1.5/2.5 = 0.6 inches,

So the scale of the map = 0.6 inches: 126 miles = 0.6 inches: 7,983,360 inches = 1: 13305600

3) line length on map = 0.76 inches, so scale = 0.76 inches: 2.1' = 0.76:(2.1*12) = 0.76:25.2 = 1:33.157

4) line length on map = 0.76 inches, so scale = 0.76: (23700:12) = 1:374210.52